# Graphical Calculus for Non-Gaussian Quantum States

### v1

@@ Abstract
Acknowledgments.— We thank Darren W. Moore for discussions. This work was supported by the National Natural Science Foundation of China (No. 12125402, No. 12350006), the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500), Beijing Natural Science Foundation (Grant No. Z240007), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K), and the Stiftung der Deutschen Wirtschaft.
Appendix A Appendix A: Proofs
@@
In this appendix, we provide the derivation details of the results presented in the main text. Since the operations $C_{e}$ commute for all hyperedges $e\in E$, the identity
@@
$\prod_{e\in E}C_{e}(t_{e})$ $=\prod_{e\in E_{a}}C_{e}(t_{e})\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})$
$=\operatorname{e}^{i\sum_{e\in E_{a}}t_{e}\operatorname{\mathit{\hat{q}}}_{e}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}\sum_{e\in\operatorname{\mathcal{A}}(a)}t_{e\cup a}\operatorname{\mathit{\hat{q}}}_{e}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})$
@@
A.1 Displacement in momentum
The momentum displacement operator $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ modifies the hyperedge set by $(E,t)\rightarrow(E,t)\cup(\{a\},s)$. That is, it adds the hyperedge $\{a\}$ with the weight $s$ if there was no such hyperedge before or changes the weight by $s$.
@@ A.1 Displacement in momentum
$=\operatorname{e}^{i(s+t_{a})\operatorname{\mathit{\hat{q}}}_{a}}\prod_{e\in E\setminus\{a\}}C_{e}(t_{e})\ket{0}^{\otimes|V|}$
@@
A.2 Displacement in position
Applying the position displacement operator $X(s)$ on the mode $a$ modifies the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ such that $(E^{\prime},t^{\prime})=(E,t)\cup(\operatorname{\mathcal{A}}(a),\tilde{t})$, where $\tilde{t}_{e}\coloneqq-s\times t_{e\cup a}$. In other words, it adds or updates the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ and multiplies the weight of the hyperedge containing $a$ by $s$:
@@ A.2 Displacement in position
$=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
@@
$=\prod_{e\in E_{a}}C_{e}(t_{e})\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\prod_{e^{\prime\prime}\in\operatorname{\mathcal{A}}(a)}C_{e^{\prime\prime}}(-s\cdot t_{e^{\prime\prime}\cup a})\operatorname{e}^{-it\operatorname{\mathit{\hat{p}}}_{a}}\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\prod_{e\in E}C_{e}(t_{e})\prod_{e^{\prime\prime}\in\operatorname{\mathcal{A}}(a)}C_{e^{\prime\prime}}(-s\cdot t_{e^{\prime\prime}\cup a})\ket{0}_{p}^{\otimes\lvert V\rvert}.$ (7)
A.3 Shearing
Applying shearing in position is trivial. This expression
@@ A.3 Shearing
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)+ish(a)\operatorname{\mathit{\hat{p}}}_{a}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
@@
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\operatorname{e}^{i\frac{s}{2}h(a)^{2}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
@@
A.4 One-mode squeezing
Applying a one-mode squeezing operation $S^{a}(s)=\operatorname{e}^{-i\frac{s}{2}(\operatorname{\mathit{\hat{p}}}_{a}\operatorname{\mathit{\hat{q}}}_{a}+\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a})}=\operatorname{e}^{-\frac{s}{2}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}}$ on the mode $a$ changes the weight of all hyperedges containing $a$ by a factor of $\operatorname{e}^{-s}$:
@@ A.4 One-mode squeezing
$=e^{-\frac{s}{2}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)\sum_{n=0}^{\infty}\frac{1}{n!}(-s)^{n}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
@@
A.5 Rotation
Applying a rotation operator in general gives us a state which is not in the form of a hypergraph state. The general form is given by:
$R^{a}(s)\ket{H}$ $=\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
@@ A.5 Rotation
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)\cos(s)}\operatorname{e}^{i\operatorname{\mathit{\hat{p}}}_{a}h(a)\sin(s)}\operatorname{e}^{i\frac{1}{4}h(a)^{2}\sin(2s)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)\cos(s)}\operatorname{e}^{i\operatorname{\mathit{\hat{p}}}_{a}h(a)\sin(s)}\operatorname{e}^{i\frac{1}{4}h(a)^{2}\sin(2s)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert-1}(R^{a}(s)\ket{0}_{p,a}),$ (12)
@@
A.6 Controlled-$Z$ gate
The effect of application of the controlled-$Z$ $C_{a,b}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{q}}}_{b}}$ gate on two modes $a,b$ directly follows from the definition of hypergraph states. If the edge $\{a,b\}\in E$ was part of the edge set, its weight gets modified, since
@@ A.6 Controlled-$Z$ gate
A.7 Measurement in $\operatorname{\mathit{\hat{q}}}$ basis
Measuring |$\outerproduct{m}{m}_{q,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$. The hyperedge set and weights get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
@@ A.7 Measurement in $\operatorname{\mathit{\hat{q}}}$ basis
$=\prod_{e\in\operatorname{\mathcal{A}}(a)}C_{e}(mt_{e\cup a})\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\outerproduct{m}{m}_{q,a}\ket{0}_{p}^{\otimes\lvert V\rvert}$
@@
A.8 Measurement in $\operatorname{\mathit{\hat{p}}}$ basis
Measuring |$\outerproduct{m}{m}_{p,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$ and creates several hyperedges in a superposition of all possible weights, with a phase determined by the weight and the measurement result $m$. The hyperedge set and weight set get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
@@ A.8 Measurement in $\operatorname{\mathit{\hat{p}}}$ basis
Figure 5: Generation of a state with $m$ hyperedges of order 3 out of a state with one hyperedge of order 3. (a) The initial hypergraph state. (b) After performing momentum shearing on mode $A$, the desired purple hyperedges of $\{B,C,D_{1}\}$,…,$\{B,C,D_{m}\}$ are generated. Additionally, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{1}}^{2}}$,…, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{m}}^{2}}$, and $e^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}}$, are produced as by-products. (c) Removing the $\{A,D_{1}\}$, …, $\{A,D_{m}\}$ edge by performing a series of two-mode controlled-$Z$ gates, $C_{A,D_{1}}(-1)$, …, $C_{A,D_{m}}(-1)$. (d) After performing the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{1}}}$, …, $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{m}}}\ket{0}_{p}^{\otimes{m+3}}$ is obtained.
Appendix B Appendix B: Further example
@@
In this section we give a further extension of the example given in the main text: In the main text in Figure 3, we couple mode $D$ with mode $A$ and use the graphical rules to show that it generates a new hyperedge of order 3 between modes $B$, $C$, and $D$. An extended example is shown in Fig. 5. We couple multiple modes $D_{1}$, …, $D_{m}$ with mode $A$ and perform the same sequence of operations which has the same effect than applying $m$ Toffoli gates on the initial state. We can see that, after performing the shearing operations on mode $A$ and a series of controlled-$Z$ gates as well as the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state Instead of applying a non-Gaussian Toffoli gate on a certain hypergraph state, we can apply a sequence of Gaussian operations on the same state which has the same effect. In the main text in Figure 3, we couple mode $D$ with mode $A$ and use the graphical rules to show that it generates a new hyperedge of order 3 between modes $B$, $C$, and $D$. An extended example is shown in Fig. 5. We couple multiple modes $D_{1}$, …, $D_{m}$ with mode $A$ and perform the same sequence of operations which has the same effect than applying $m$ Toffoli gates on the initial state. We can see that, after performing a series of shearing operations and controlled-$Z$ gates, the new hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{1}}},\dots,\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{m}}}\ket{0}_{p}^{\otimes{m+3}}$ is obtained.

### v2

@@ 4.3 Acknowledgments
We thank Darren W. Moore for discussions. This work was supported by the National Natural Science Foundation of China (No. 12125402, No. 12350006), the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500), Beijing Natural Science Foundation (Grant No. Z240007), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K), and the Stiftung der Deutschen Wirtschaft.
Appendix A Derivations of graphical calculus
@@
In the following, we provide the derivation details of the results presented above. Since the operations $C_{e}$ commute for all hyperedges $e\in E$, the identity
@@
$\prod_{e\in E}C_{e}(t_{e})$ $=\operatorname{e}^{i\sum_{e\in E_{a}}t_{e}\operatorname{\mathit{\hat{q}}}_{e}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})$
@@
Displacement in momentum: The momentum displacement operator $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ modifies the hyperedge set by $(E,t)\rightarrow(E,t)\cup(\{a\},s)$. That is, it adds the hyperedge $\{a\}$ with the weight $s$ if there was no such hyperedge before or changes the weight by $s$.
@@
Displacement in position: Applying the position displacement operator $X(s)$ on the mode $a$ modifies the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ such that $(E^{\prime},t^{\prime})=(E,t)\cup(\operatorname{\mathcal{A}}(a),\tilde{t})$, where $\tilde{t}_{e}\coloneqq-s\times t_{e\cup a}$. In other words, it adds or updates the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ and multiplies the weight of the hyperedge containing $a$ by $s$:
@@
$=\prod_{e\in E}C_{e}(t_{e})\prod_{e^{\prime\prime}\in\operatorname{\mathcal{A}}(a)}C_{e^{\prime\prime}}(-s\dot{t}_{e^{\prime\prime}\cup a})\ket{0}_{p}^{\otimes\lvert V\rvert}.$ (5)
Shearing: Applying shearing in position is trivial. This expression
@@
One-mode squeezing: Applying a one-mode squeezing operation $S^{a}(s)=\operatorname{e}^{-i\frac{s}{2}(\operatorname{\mathit{\hat{p}}}_{a}\operatorname{\mathit{\hat{q}}}_{a}+\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a})}=\operatorname{e}^{-\frac{s}{2}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}}$ on the mode $a$ changes the weight of all hyperedges containing $a$ by a factor of $\operatorname{e}^{-s}$:
@@
Rotation: Applying a rotation operator in general gives us a state which is not in the form of a hypergraph state. The general form is given by:
$R^{a}(s)$ |$\ket{H}=\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
@@
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)\cos(s)}\operatorname{e}^{i\operatorname{\mathit{\hat{p}}}_{a}h(a)\sin(s)}\operatorname{e}^{i\frac{1}{4}h(a)^{2}\sin(2s)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}\ket{0}_{p}^{\otimes\lvert V\rvert}$ (10)
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)\cos(s)}\operatorname{e}^{i\operatorname{\mathit{\hat{p}}}_{a}h(a)\sin(s)}\operatorname{e}^{i\frac{1}{4}h(a)^{2}\sin(2s)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert-1}(R^{a}(s)\ket{0}_{p,a}),$
@@
Controlled-$Z$ gate: The effect of application of the controlled-$Z$ $C_{a,b}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{q}}}_{b}}$ gate on two modes $a,b$ directly follows from the definition of hypergraph states. If the edge $\{a,b\}\in E$ was part of the edge set, its weight gets modified, since
@@
Measurement in $\operatorname{\mathit{\hat{q}}}$ basis: Measuring |$\outerproduct{m}{m}_{q,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$. The hyperedge set and weights get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
@@
Measurement in $\operatorname{\mathit{\hat{p}}}$ basis: Measuring |$\outerproduct{m}{m}_{p,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$ and creates several hyperedges in a superposition of all possible weights, with a phase determined by the weight and the measurement result $m$. The hyperedge set and weight set get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
@@
Appendix B Effects of Finite Squeezing
According to the hypergraph state definition defined before, every vertex represents a infinitely squeezed zero-momentum eigenstate |$\ket{0}_{p}=\lim_{r\rightarrow\infty}S(-r)\ket{0}$, where |$\ket{0}$ is the vacuum state. In experimental realisations, these states have to be approximated by finitely squeezed states |$\ket{0}_{p}\approx S(-r)\ket{0}$ for some large $r\in\mathbb{R}$. The hypergraph state |$\ket{H_{\text{ideal}}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\ket{000}_{p}$ corresponding to the three vertex hypergraph $H$ with a single hyperedge $\{A,B,C\}$ is therefore approximated by |$\ket{H_{\text{approx}}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}S_{A}(-r_{A})S_{B}(-r_{B})S_{C}(-r_{C})\ket{000}$, where $r_{A},r_{B},r_{C}\in\mathbb{R}$. In the derivation of the graphical rules, we make use of the identity $\operatorname{e}^{if(\operatorname{\mathit{\hat{q}}},\operatorname{\mathit{\hat{p}}})\operatorname{\mathit{\hat{p}}}}\ket{0}_{p}=\ket{0}_{p}$ which holds for all functions $f(\operatorname{\mathit{\hat{q}}},\operatorname{\mathit{\hat{p}}})$ of $\operatorname{\mathit{\hat{q}}}$ and $\operatorname{\mathit{\hat{p}}}$. However, in the case of finite squeezing, terms of the form $\operatorname{e}^{if(\operatorname{\mathit{\hat{q}}},\operatorname{\mathit{\hat{p}}})\operatorname{\mathit{\hat{p}}}}$ do not vanish. For example, applying the position displacement operator $X^{A}(s)$ on the two states |$\ket{H_{\text{ideal}}}$ and |$\ket{H_{\text{approx}}}$ leads to
$X^{A}(s)\ket{H_{\text{ideal}}}=$ $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\ket{000}_{p},$
$X^{A}(s)\ket{H_{\text{approx}}}=$ $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-is\hat{p}_{A}}S_{A}(-r_{A})S_{B}(-r_{B})S_{C}(-r_{C})\ket{000}.$ (16)
We can predict the state |$\ket{H_{\text{ideal}}}$ by
|$\ket{H_{\text{predict}}}=$ $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}S_{A}(-r_{A})S_{B}(-r_{B})S_{C}(-r_{C})\ket{000}.$ (17)
The fidelity of |$\ket{H_{\text{predict}}}$ and the actually displaced state $X^{A}(s)\ket{H_{\text{approx}}}$ is given by $\operatorname{e}^{-\frac{1}{2}s^{2}\operatorname{e}^{-2r_{A}}}$, which tends to 1 for a large squeezing parameter $r_{A}$. Here, we assume that the states can be squeezed such that the fidelity between the actually prepared state and the state computed using the graphical rules from this paper is sufficiently high. Other rules can be calculated fidelity in a similar way.
@@ Appendix B Effects of Finite Squeezing
Figure 5: Generation of a state with $m$ 3rd-order hyperedges. (a) The initial hypergraph state with only one 3rd-order hyperedge. (b) After performing momentum shearing on mode $A$, the desired purple hyperedges of $\{B,C,D_{1}\}$,…,$\{B,C,D_{m}\}$ are generated. Additionally, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{1}}^{2}}$,…, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{m}}^{2}}$, and $e^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}}$, are produced as by-products. (c) Removing the $\{A,D_{1}\}$, …, $\{A,D_{m}\}$ edge by performing a series of two-mode controlled-$Z$ gates, $C_{A,D_{1}}(-1)$, …, $C_{A,D_{m}}(-1)$. (d) After performing the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{1}}}$$\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{m}}}\ket{0}_{p}^{\otimes{m+3}}$ is obtained.
Appendix C Further Extension of the Example Given in Fig. 3
@@
In Fig. 3, we couple mode $D$ with mode $A$ and use the graphical rules to show that it generates a new hyperedge of order 3 between modes $B$, $C$, and $D$. An extended example is shown in Fig. 5. We couple multiple modes $D_{1}$, …, $D_{m}$ with mode $A$ and perform the same sequence of operations which has the same effect than applying $m$ Toffoli gates on the initial state.
We can see that, after performing the shearing operations on mode $A$ and a series of controlled-$Z$ gates as well as the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state Instead of applying a non-Gaussian Toffoli gate on a certain hypergraph state, we can apply a sequence of Gaussian operations on the same state which has the same effect.